A particle on a stretched string supporting a travelling wave, takes 5.0 ms to move from its mean position to the extreme position. The distance between two consecutive particles, which are at their mean positions, is 2.0 cm. Find the frequency, the wavelength and the wave speed.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. Time period, T = 4 * 5 ms = 20 * 10–3 = 2 * 10–2 s λ = 2 * 2 cm = 4 cm frequency, f = 1/T = 1/(2 * 10–2) = 50 s–1 = 50 Hz Wave speed = λ f = 4 * 50 m/s = 2000 m/s = 2 m/s